OpenStudy (anonymous):

A parabola has a vertex at (4, 0) and passes through the point (6, 1). Which of the following is the equation of this parabola in the form 2003-18-03-00-00_files/i0230000.jpg?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

The following work represents an attempt to find the equation of the parabola which has a vertex of (0, 0) and passes through the point (1, 3). Is the work correct? If not, what is the error? 2003-18-03-00-00_files/i0240000.jpg

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

and fan

OpenStudy (anonymous):

Well, if the vertex is at (4,0), we can plug in to the standard form like this\[y=a(x-h)^2+k=a(x-4)^2+0=a(x-4)^2\]Now substitute (x,y)=(6,1) to solve for a.

OpenStudy (anonymous):

That goes like this\[1=a(6-4)^2=4a \implies a=\frac{1}{4}\]

OpenStudy (anonymous):

That, in turn, means your function is\[y=\frac{1}{4}(x-4)^2=\frac{x^2}{4}-2x+4\]

OpenStudy (anonymous):

Do math every day.